OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

m=-2

OpenStudy (anonymous):

no no m=2

OpenStudy (anonymous):

can you help me in part 2 please

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

y=2x-2

OpenStudy (anonymous):

(3, 4) and (–3, –8) in point-slope form. m = (4 + 8) / (3 + 3) = 12/6 = 2 y - 4 = 2 (x -3) =>y = 2x -2

OpenStudy (anonymous):

doesn't matter which value you use for your point. Both will work out the same. they are both on the same line, that is why it does not matter which one you pick.

OpenStudy (anonymous):

can some one please help me on PART 2 pleaseee

OpenStudy (anonymous):

Use x , y intercept points are the most convenient ways to graph

OpenStudy (anonymous):

i picked the point (3,4) cus it doesn't have negative values and that's eaiser :p